A group of dogs consists of six brown males, two brown females, six white males, four white females, four black males, and six black females.

Q1: What is the probability of selecting at random a brown female dog?

Q2: What is the conditional probability of selecting at random a black female dog and then a male white dog?

Q3: What is the conditional probability of selecting at random a brown dog, if the dog is a female?
Answer #1

Q1: Required probability = 2/(6+2+6+4+4+6) = 0.071

Q2: Required probability = (6/28)*(6/27) = 0.048

Q3: Required probability = 2/(2+4+6) = 0.167

Please give me a thumbs-up if this helps you out. Thank you!